If f(x) = [x], where [x] denotes the integral part of x, then f(x) is continuous at all integral values of x ?

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
f(x) = [x]
f(x) = 0, 0\leq x< 1
= 1, 1\leq x< 2
= 2, 2\leq x< 3
= 3, 3\leq x< 4
Similarly for all the other integer values, so f(x) is discontinuous at all the integer values.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty